This is about parabola's.
I know the fomula to find the vertex is y=aX^2+bx+c
But I dont understand the process of working the problem out.
The equation is y=(x-2)^2+3
(I dont have the raised to 2nd power key so I put ^2 hope its not a problem) Answer by lwsshak3(11628) 
 This is about parabola's.
I know the fomula to find the vertex is y=aX^2+bx+c
But I dont understand the process of working the problem out.
The equation is y=(x-2)^2+3
**
Standard form of equation for a parabola which opens upwards: y=A(x-h)^2=k, with (h,k) being the (x,y) coordinates of the vertex.
Given equation is already in standard form from which you can read the vertex.
Vertex: (2,3)